If the quadratic equation ax²-bx+12=0,where a and B are positive integers not exceeding 10,has roots both greater than 2,then number of possible ordered pair (a, b) is

Grade:11

1 Answers

Lucky
6 years ago
as alpha >2 and beta >2 . So alpha+ beta >4 and Alpha . Beta >4 . So alpha + beta =b/a. and Alpha . Beta = 12/a By putting these and solving for a we get a8 but b0 So a=1,2. Now when a=1weget b>4 ,but we know that b